### v1

@@ Abstract
To evaluate the performance of quantum computing systems relative to classical counterparts and explore the potential for quantum advantage, we propose a game-solving benchmark based on Elo ratings in the game of tic-tac-toe. We compare classical convolutional neural networks (CNNs), quantum convolutional neural networks (QCNNs), and hybrid classical-quantum models by assessing their performance against a random-move agent in automated matches. Additionally, we implement a QCNN integrated with quantum communication and evaluate its performance to quantify the overhead introduced by noisy quantum channels. Our results show that the classical-quantum hybrid model achieves Elo ratings comparable to those of classical CNNs, while the standalone QCNN underperforms under current hardware constraints. The communication overhead was found to be modest. These findings demonstrate the viability of using game-based benchmarks for evaluating quantum computing systems and suggest that quantum communication can be incorporated with limited impact on performance, providing a foundation for future hybrid quantum applications.
@@
Quantum Machine Learning, Quantum Communication, Quantum/Classical Algorithm, Rating System, Tic-Tac-Toe Solver
@@ I Introduction
Quantum computers have brought new possibilities to many previously difficult problems for classical computers to solve [1, 2]. In recent years, the development of quantum computers has made remarkable progress. Noise Intermediate-Scale Quantum (NISQ) computers exceeding 1,000 qubits have emerged, and it is anticipated that early Fault-Tolerant Quantum Computers (FTQC) will also appear in the not-so-distant future [3]. With further advancements in both hardware and algorithms, the potential for applying quantum advantage to real-world problems is increasingly becoming a reality.
Current benchmarks for quantum computers, such as random circuit sampling, demonstrate quantum advantages in limited tasks [4]. However, evidence supporting their superiority in solving real-world problems remains inconclusive [5]. Furthermore, the absence of common and standardized metrics for computational performance presents a significant challenge in evaluating the potential societal and practical impact of future quantum computing technologies.
Classical game AIs, such as AlphaZero [6], have been highly successful and have significantly outperformed human abilities in perfect-information zero-sum games. While quantum game AIs have been proposed for solving single-player puzzle games using search algorithms [7], algorithms that can adapt to dynamically changing board information in competitive games remain unexplored. By utilizing a unified metric, it becomes possible to compare the performance of quantum and classical systems directly.
Figure 1: The game solver initially runs on the client, receiving the board state and selecting the next move using a classical, quantum, or hybrid neural network. In the Quantum Internet setup, a noisy quantum communication channel connects the client and server. The client processes the input with a classical neural network, encodes the output into qubits, and sends them through the quantum channel. The server then applies a quantum neural network and returns either classical measurement results or unmeasured qubits. The client uses the received data to determine the next move.
@@
The objective of this study is to quantitatively assess the capabilities of quantum computers using competitive games. To this end, we adopt the following methodology:
@@
    Design four policy types—random, classical, quantum, and classical-quantum hybrid—and compare their performance.
@@
Furthermore, we evaluate the impact of connecting quantum devices with varying performance levels, in anticipation of future quantum communication applications. Just as classical high-performance computing is widely accessed via the cloud, similar usage scenarios are expected for quantum devices. In this study, we examine the practicality of quantum applications over the quantum Internet (QI) by analyzing the communication overhead caused by noise, assuming a client with limited quantum resources and a server with abundant ones. While the detailed QI protocols involving entanglement are beyond our scope, we construct and test simplified models under two scenarios, based on the client’s quantum resource availability.
To explore this approach, we first review the machine learning techniques and Elo rating system used in our study (Sec. II). We then describe the architecture of the proposed game engines (Sec. III) and evaluate their performance (Sec. IV). Next, we introduce the quantum Internet model and assess the performance of game engines operating within this setting (Sec. V). Finally, we conclude with a summary of our findings and discuss future research directions (Sec. VI).
@@ II Preliminaries
Reinforcement learning is a method of learning to make the optimal choice in a particular environment. Unlike supervised and unsupervised learning, it aims to learn through the interaction between the environment and the agent and maximizing the reward so that the agent can cope with unknown domains.
@@
There is a strong connection between reinforcement learning and game AI, especially in games such as chess and shogi. Competitive games are well-suited for reinforcement learning because outcomes (win/loss) are clearly defined and player performance can be quantitatively evaluated. As reinforcement learning algorithms have evolved, game AI has improved in parallel.
Q-learning is one of the foundational algorithms in reinforcement learning. In this study, we developed and evaluated a game engine based on Q-learning. This section provides a brief overview of the algorithm [8].
@@
$Q(s_{i},a_{i})\leftarrow Q(s_{i},a_{i})\\ +\alpha\Bigl{[}r_{i+1}+\gamma\max_{a_{i+1}}Q(s_{i+1},a_{i+1})-Q(s_{i},a_{i})\Bigr{]}$ (1)
@@ II-B Quantum Machine Learning
Quantum machine learning [9, 10] has emerged as a promising application of quantum computing. It aims to develop learning algorithms that leverage quantum properties to outperform classical machine learning in terms of computational efficiency and model expressiveness.
In particular, quantum neural networks (QNNs) have been proposed as quantum analogues of classical neural networks. When processing classical data, a typical QNN pipeline involves three stages: encoding the data into quantum states, processing it using parameterized quantum circuits (ansatz), and measuring the output to obtain classical values. These values are used to optimize parameters via classical computation.
@@
One such approach is the hybrid neural network (HNN), which combines classical and quantum components. For example, classical neural networks (CNNs) may be placed before or after a quantum layer, or quantum circuits may be used to perform certain computations such as gradient descent [11, 12]. These hybrid models aim to exploit the strengths of both paradigms, making them suitable for near-term quantum devices.
Another notable architecture is the quantum convolutional neural network (QCNN) [13], which is inspired by classical convolutional networks but designed to operate on quantum states. QCNNs can achieve more efficient learning with fewer parameters than general QNNs. Moreover, they have been shown to mitigate the barren plateau problem [14], in which gradients vanish during training, hindering effective optimization.
The QCNN ansatz consists of alternating convolution and pooling layers (Fig. 16). In the convolutional layers, entanglement between qubits is used to extract local features. The pooling layers reduce dimensionality by discarding a subset of qubits, rather than aggregating information as in classical CNNs. Repeating these layers enables hierarchical feature extraction and supports scalable learning.
These architectures demonstrate how quantum machine learning can be structured to balance expressivity and trainability. Hybrid models are particularly promising in the current era of noisy intermediate-scale quantum (NISQ) devices, while QCNNs offer a scalable framework for leveraging quantum entanglement and feature abstraction in learning tasks.
@@
Elo rating [15, 16, 17] is a widely used system for quantifying the strength of individual players in competitive games, such as chess. It enables an approximate comparison of player skill levels, regardless of game order or the number of opponents.
@@ II-D1 Overview of the Client-Server Model
Quantum computers require large-scale infrastructure, so they are typically owned by major vendors rather than individual users. In this situation, the client sends some classical or quantum information to the server owned by the vendors, and then the vendor executes the costly part of the computation. This type of computational setup is referred to as the client-server model, and it is regarded as an application of the quantum Internet. The client-server model not only enables physically distributed computing systems but also offers a variety of additional advantages when quantum machine learning tasks use this model[18, 19]. One can consider a scenario where the client-server model is adopted in the following manner: the client encodes classical data into quantum states and sends them to the server; the server calculate ansatz to perform the necessary quantum operations; then, the resulting quantum states are returned to the client, who performs the measurements and processes the outcomes.
@@ III Tic-Tac-Toe Engines
In this section, we describe three types of game engines, each utilizing a classical, quantum, or hybrid neural network architecture. All neural networks are trained using classical optimization methods. Quantum circuits are integrated into the PyTorch framework via the TorchConnector provided by Qiskit [20]. We employ Huber Loss [21] as the loss function and the Adam optimizer [22] for weight updates.
@@ III-A Classical neural networks (CNN, CCNN)
Figure 7: Data flow in the classical convolutional neural network (CCNN) engine. The tic-tac-toe board is encoded into numerical values and passed through convolutional layers, followed by fully connected layers. The output consists of nine values corresponding to the board cells, and the move with the highest score is selected.
III-B Quantum convolutional neural networks (QCNN)
The QCNN-based engines differ in their choice of embedding circuits (HEE, TPE, ZFeatureMap, ZZFeatureMap). All models use 18 qubits, with the 9-cell input duplicated to match the quantum register size. Details of the quantum circuit designs are provided in Appendix A-A.
The number of shots in the measurement is set to 1000, and half of each qubit is measured in the Z basis, and the expected value is taken as the output.
@@ III-B Quantum convolutional neural networks (QCNN)
III-C Hybrid neural networks (CQCNN)
The hybrid engine (CQCNN) combines classical and quantum neural network components. The input and output layers are implemented using classical convolutional neural networks (CNNs), while the hidden layer consists of a quantum neural network, either a QNN or a QCNN, depending on the configuration (Fig. 9).
To accommodate quantum processing, the input board state is first converted into a 9-dimensional vector using the same encoding as in the classical models (O: +1, X: –1, empty: 0). This vector is then expanded to match the number of qubits used in the quantum layer (7 or 18 qubits, depending on the setup), and embedded into quantum states using one of the four encoding methods: HEE, TPE, ZFeatureMap, or ZZFeatureMap.
@@ III-C Hybrid neural networks (CQCNN)
The quantum layer is followed by measurement in the Z-basis. We consider two measurement strategies:
@@
In the QNN-based configurations, two types of parameterized quantum circuits (ansatz) are used: EfficientSU2 and RealAmplitudes, both provided by Qiskit. In contrast, the QCNN-based models use a fixed QCNN ansatz with a convolution–pooling structure, and do not incorporate these two ansatz types.
@@
Figure 9: Data flow in the hybrid classical–quantum–classical neural network (CQCNN) engine. Encoded board data is processed through an input CNN layer, followed by a quantum layer (QNN or QCNN), and a final output CNN layer. The resulting nine output values are used to determine the next move.
The number of shots in the measurement is set to 1000. We output from the quantum layer in 2 different methods, Estimator and Sampler. Estimator outputs the expected value for each observable after the measurement. Sampler outputs the quasi-probabilities after the measurement. In these engines we used, qubits are measured in the Z basis.
@@ IV Evaluation of Engines
In this section, we evaluate 46 game engine configurations across classical, quantum, and hybrid neural network architectures. Each engine is trained using the reinforcement learning method described in Section II and evaluated using Elo ratings obtained through self-play against a random policy.
@@ IV-A Overview of Evaluation Settings
The 46 engines are categorized as follows:
@@
    QCNN engines: 4 configurations using different embedding methods (Table I (b)).
@@
    Hybrid CQCNN engines (QNN layer): 32 variants (2 measurement methods × 2 qubit counts × 8 configurations).
@@
    Hybrid CQCNN engines (QCNN layer): 8 variants with 2 qubit settings and 4 embedding types.
TABLE I: NN Architectures: (a) CNN: Stronger has 3 layers, Weaker has 2, both using $3\times 3$ kernels. Stronger also has larger fully connected layers. (b) QCNN: A simple architecture with a single QCNN layer, where the input board is duplicated to match its inputs. (c) CQCNN: $N=7$ or $18$ for QNN, $10$ or $18$ for QCNN. CQCNNs with a QNN hidden layer support Sampler and Estimator outputs, while those with a QCNN hidden layer support only Estimator.
@@
Fully Connected - $9$
(b) QCNN Architectures
@@
Sampler Fully Connected 1 $N$
@@
Fully Connected 2 $9$
Estimator Fully Connected 1 $N$
QNN/QCNN $N$
Fully Connected 2 $9$
These engines used the Tanh function as the activation function before outputting. All engines were trained for 250 games and evaluated over 10,000 games. Elo ratings were updated every 100 games using $K=1$ in Eq. (3).

### v2

@@ Abstract
To evaluate the performance of quantum computing systems relative to classical counterparts and explore the potential, we propose a game-solving benchmark based on Elo ratings in the game of tic-tac-toe. We compare classical convolutional neural networks (CCNNs), quantum or quantum convolutional neural networks (QNNs, QCNNs), and hybrid classical-quantum neural networks (Hybrid NNs) by assessing their performance based on round-robin matches. Our results show that the Hybrid NNs engines achieve Elo ratings comparable to those of CCNNs engines, while the quantum engines underperform under current hardware constraints. Additionally, we implement a QNN integrated with quantum communication and evaluate its performance to quantify the overhead introduced by noisy quantum channels, and the communication overhead was found to be modest. These results demonstrate the viability of using game-based benchmarks for evaluating quantum computing systems and suggest that quantum communication can be incorporated with limited impact on performance, providing a foundation for future hybrid quantum applications.
@@
Quantum Machine Learning, Quantum Communication, Quantum/Classical Algorithm, Quantum Advantage, Rating System, Tic-Tac-Toe Solver
@@ I Introduction
Quantum computers have brought new possibilities to many previously difficult problems for classical computers to solve [Shor_1997, grover1996fast]. Furthermore, the emergence of an algorithm that allows many existing algorithms to be viewed from a unified perspective is also accelerating the development of new quantum algorithms [PRXQuantum.2.040203, 10.1145/3549524]. In recent years, the development of quantum computers has made remarkable progress. Noisy Intermediate-Scale Quantum (NISQ) computers exceeding 1,000 qubits have emerged, and research and development for Fault-Tolerant Quantum Computers (FTQC) has become very active [yoder2025tourgrossmodularquantum]. With further advancements in both algorithms and hardware, the potential for applying quantum advantage to real-world problems is increasingly becoming a reality.
Current benchmarks for quantum computers, such as random circuit sampling, demonstrate quantum advantages in limited tasks [morvan2024phase]. However, evidence supporting their superiority in solving real-world problems remains inconclusive [kim2023evidence]. Furthermore, the absence of common and standardized metrics for computational performance presents a significant challenge in evaluating the potential societal and practical impact of future quantum computing technologies. Therefore, it is necessary to measure the performance of quantum systems from multiple angles using various benchmarking methods, such as Quantum Volume [cross2019validating] and Circuit Mirroring [proctor2022measuring], the latter being a flexible technique capable of generating customized, efficiently verifiable benchmarks based on any circuit structure, which can even transform Quantum Volume into a fully scalable measure.
Classical game AIs, such as AlphaZero [doi:10.1126/science.aar6404], have been highly successful and have significantly outperformed human abilities in perfect-information zero-sum games. While quantum game AIs have been proposed for solving single-player puzzle games using search algorithms [IBMQchallenge2020], algorithms that can adapt to dynamically changing board information in competitive games remain unexplored.
By utilizing a unified metric, it becomes possible to compare the performance of quantum and classical systems directly. Being able to compare these systems in a unified environment allows us to evaluate the progress of quantum systems not only from a quantum perspective but also from a classical one. This has the potential to serve as a key benchmark for demonstrating quantum advantage.
Figure 1: The game solver initially runs on the client, receiving the board state and selecting the next move using a classical, quantum, or hybrid neural network. In the Quantum Internet setup, a noisy quantum communication (Q-comm) channel connects the client and server. The client processes the input with a classical neural network, encodes the output into qubits, and sends them through the quantum channel. The server then applies a quantum neural network and returns either classical measurement results or unmeasured qubits. The client uses the received data to determine the next move.
@@
The objective of this research is to quantitatively assess the capabilities of quantum computers using competitive games. To this end, we adopt the following methodology:
@@
    Design three policy types—classical, quantum, and classical-quantum hybrid—and compare their performance.
@@
Furthermore, we evaluate the impact of connecting quantum devices with varying performance levels, in anticipation of future quantum communication applications. Just as classical high-performance computing is widely accessed via the cloud, similar usage scenarios are expected for quantum devices. In this research, we examine the practicality of quantum applications over the quantum Internet (QI) by analyzing the communication overhead caused by noise, assuming a client with limited quantum resources and a server with abundant ones. While the detailed QI protocols involving entanglement are beyond our scope, we construct and test simplified models under two scenarios, based on the client’s quantum resource availability.
To explore this approach, we first review the machine learning techniques and Elo rating system used in our work (Sec. II). We then describe the architecture of the proposed game engines (Sec. III) and evaluate their performance (Sec. IV). Next, we introduce the QI model and assess the performance of game engines operating within this setting (Sec. V). Finally, we give a summary and discuss future research directions (Sec. VI).
@@ II Preliminaries
Reinforcement learning [sutton1998reinforcement, watkins1989learning, mnih2013playing] is a method of learning to make the optimal choice in a particular environment. Unlike supervised and unsupervised learning, it aims to learn through the interaction between the environment and the agent and maximize the reward so that the agent can cope with unknown domains.
@@
There is a strong connection between reinforcement learning and game AI, especially in games such as chess and shogi [doi:10.1126/science.aar6404, Schrittwieser_2020]. Competitive games are well-suited for reinforcement learning because outcomes (win/loss) are clearly defined and player performance can be quantitatively evaluated. As reinforcement learning algorithms have evolved, game AI has improved in parallel.
Q-learning [watkins1992q] is one of the foundational algorithms in reinforcement learning. In this research, we developed and evaluated a game engine based on Q-learning. This section provides a brief overview of the algorithm.
@@
$Q(s_{i},a_{i})\leftarrow Q(s_{i},a_{i})\\ +\alpha\Bigl[r_{i+1}+\gamma\max_{a_{i+1}}Q(s_{i+1},a_{i+1})-Q(s_{i},a_{i})\Bigr].$ (1)
@@ II-B Quantum Machine Learning
Quantum machine learning [biamonte2017quantum, schuld2015introduction] has emerged as a promising application of quantum computing. It aims to develop learning algorithms that leverage quantum properties to outperform classical machine learning in terms of computational efficiency and model expressiveness.
In particular, quantum neural networks (QNNs) [gupta2001quantum, beer2020training] have been proposed as quantum analogues of classical neural networks. When processing classical data, a typical QNN pipeline involves three stages: encoding the data into quantum states, processing it using parameterized quantum circuits (ansatz), and measuring the output to obtain classical values. These values are used to optimize parameters via classical computation.
@@
One such approach is the hybrid neural network (HNN), which combines classical and quantum components. For example, classical neural networks (CNNs) may be placed before or after a quantum layer, or quantum circuits may be used to perform certain computations such as gradient descent [kerenidis2020quantum, stokes2020quantum]. These hybrid models aim to exploit the strengths of both paradigms, making them suitable for near-term quantum devices.
Another notable architecture is the quantum convolutional neural network (QCNN) [cong2019quantum], which is inspired by classical convolutional networks but designed to operate on quantum states. Compared to general QNNs, QCNNs can learn more efficiently with fewer parameters, as the number of parameters scales logarithmically with the number of qubits. Moreover, they have been shown to mitigate the barren plateau problem [pesah2021absence], in which gradients vanish during training, hindering effective optimization.
The QCNN ansatz consists of alternating convolution and pooling layers (Fig. 15). In the convolutional layers, entanglement between qubits is used to extract local features. While the pooling layers in classical CNNs aggregate information, its quantum counterparts achieve dimensionality reduction by concentrating information onto a single qubit, which is then discarded. Repeating these layers enables hierarchical feature extraction and supports scalable learning.
These architectures demonstrate how quantum machine learning can be structured to balance expressivity and trainability. Hybrid models are particularly promising in the current era of NISQ devices, while QCNNs offer a scalable framework for leveraging quantum entanglement and feature abstraction in learning tasks.
@@
Elo rating [elo1978, albers2001elo, HVATTUM2010460] is a widely used system for quantifying the strength of individual players in competitive games, such as chess. It enables an approximate comparison of player skill levels, regardless of game order or the number of opponents.
@@ II-D1 Overview of the Client-Server Model
Quantum computers require large-scale infrastructure, so they are typically owned by major vendors rather than individual users. In this situation, the client sends some classical or quantum information to the server owned by the vendors, and then the vendor executes the costly part of the computation. This type of computational setup is referred to as the client-server model, and it is regarded as an application of the quantum Internet. The client-server model not only enables physically distributed computing systems but also offers a variety of additional advantages when quantum machine learning tasks use this model[araujo2024quantum, song2024quantum]. One can consider a scenario where the client-server model is adopted in the following manner: the client encodes classical data into quantum states and sends them to the server; the server calculate ansatz to perform the necessary quantum operations; then, the resulting quantum states are returned to the client, who performs the measurements and processes the outcomes.
@@ III Tic-Tac-Toe Engines
In this section, we describe three types of game engines, each utilizing a classical, quantum, or hybrid neural network architecture. All neural networks are trained using classical optimization methods. Quantum circuits are integrated into the PyTorch framework via the TorchConnector provided by Qiskit [qiskit2024]. We employ Huber Loss [huber1992robust] as the loss function and the Adam optimizer [kingma2014adam] for weight updates.
@@ III-A Classical neural networks (CNN, CCNN)
Figure 7: Data flow in the classical convolutional neural network (CCNN) engine. The tic-tac-toe board is encoded into numerical values and passed through convolutional layers, followed by fully connected (FC) layers. The output consists of nine values corresponding to the board cells, and the move with the highest score is selected.
III-B Quantum neural networks (QNN, QCNN)
The QNN-based engines differ in their choice of embedding circuits (ZFeatureMap, ZZFeatureMap, HEE, TPE), and ansatz circuits (RealAmplitudes, EfficientSU2). All models use 9 qubits, matching the size of board.
The QCNN-based engines (Fig. 8) differ in their choice of embedding circuits mensioned above, and ansatz circuit is QCNN. All models use 18 qubits, with the 9-cell iput duplicated to match the quantum register size. Details of the quantum circuit designs are provided in Appendix A.
The number of shots in the measurement is set to 1024, and half of each qubit is measured in the Z basis, and the expected value is taken as the output.
@@ III-B Quantum neural networks (QNN, QCNN)
III-C Hybrid neural networks (HNN)
The hybrid engine combines classical and quantum neural network components. The input and output layers are implemented using classical convolutional neural networks (CNNs), while the hidden layer consists of a quantum neural network, either a QNN or a QCNN, depending on the configuration (Fig. 9). We refer to this architecture as HNN, based on the order in which the classical and quantum neural network components are combined.
To accommodate quantum processing, the input board state is first converted into a 9-dimensional vector using the same encoding as in the classical models (O: +1, X: –1, empty: 0). This vector is then expanded to match the number of qubits used in the quantum layer (8 or 16 qubits, depending on the setup), and embedded into quantum states using one of the four encoding methods: ZFeatureMap, ZZFeatureMap, HEE, or TPE.
@@ III-C Hybrid neural networks (HNN)
The quantum layer is followed by measurement in the Z-basis and the number of shots in the measurement is set to 1024. We consider two measurement strategies:
@@
In the QNN-based configurations, two types of parameterized quantum circuits (ansatz) are used: EfficientSU2 and RealAmplitudes, both provided by Qiskit. In contrast, the QCNN-based models use a fixed QCNN ansatz with a convolution–pooling structure, and do not incorporate these two ansatze types.
@@
Figure 9: Data flow in the hybrid classical-quantum neural network (HNN) engine. Encoded board data is processed through an input CNN layer, followed by a quantum layer (QNN or QCNN), and a final output CNN layer. The resulting nine output values are used to determine the next move.
We output from the quantum layer in two different methods, Estimator and Sampler as mensioned above. Estimator outputs the expected value for each observable after the measurement. Sampler outputs the quasi-probabilities after the measurement. In these engines we used, qubits are measured in the Z basis.
@@ IV Evaluation of Engines
In this section, we evaluate 54 game engine configurations across classical, quantum, and hybrid neural network architectures. Each engine is trained using the reinforcement learning method described in Section II and evaluated using Elo ratings based on round-robin matches.
@@ IV-A Overview of Evaluation Settings
The 54 engines are categorized as follows:
@@
    QNN engines: 8 configurations using different embedding and ansatz circuits (Table I (b)).
@@
    QCNN engines: 4 configurations using different embedding types (Table I (c)).
@@
    Hybrid NN engines (QNN layer): 32 variants (2 measurement methods × 2 qubit counts × 8 configurations).
  5. 5.
    Hybrid NN engines (QCNN layer): 8 variants with 2 qubit settings and 4 embedding types.
TABLE I: NN Architectures: (a) CNN: Stronger has 3 layers, Weaker has 2, both using $3\times 3$ kernels. Stronger also has larger fully connected layers. (b) QNN: A single QNN layer using 9 qubits. (c) QCNN: A single QCNN layer, where the input board is duplicated to match its inputs. (d) HNN: $N=8$ or $16$ for QNN, 88 or $16$ for QCNN. HNNs with a QNN hidden layer support Sampler and Estimator outputs, while those with a QCNN hidden layer support only Estimator.
@@
Fully Connected - 99
(b) QNN Architectures
Layer Output Size
Flatten 99
QNN 99
(c) QCNN Architectures
@@
Sampler Fully Connected 1 NN
@@
Fully Connected 2 99
Estimator Fully Connected 1 NN
QNN/QCNN NN
Fully Connected 2 99
These engines used the Tanh function as the activation function before outputting.
We adopt epsilon-greedy decay strategy to train all engines by self-play which is using same engines for battle to update weight parameters. The evaluation was conducted based on round-robin matches, battled for 100 games per one engine. Elo ratings were updated every 100 games from initial rate 1500, using $K=32$ in Eq. (3).